OpenStudy (anonymous):

Write a compound probability problem that uses dependent events involving a standard deck of cards. Solve your problem.

OpenStudy (anonymous):

Dependent: the outcome of events is affected by previous events What is the probability that you will draw two clubs in a row from a standard deck First card you draw has to be a club, there are 13 clubs & 52 cards, so probability is 13/52 Second card you draw has to be a club, there are now 12 clubs & 51 cards, so probability 12/51. Here the first event, pulling a card from the deck, changed the probability for this second event, so they are dependent. 13/52 * 12/51 = 156/2652 = 1/17 ******** Independent means that each event is NOT affected by previous events What is the probability that the first heart you draw will be a queen & the first diamond that you draw will be a queen Hearts: 1 queen in 13 hearts, so probability 1/13 Diamonds: 1 queen in 13 diamonds (see, doesn't matter what your first heart was, so independent), so probability 1/13 1/13 * 1/13 = 1/169 http://answers.yahoo.com/question/index?qid=20130225210850AAQT8dL

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

Anytime